Plan:

Start with LBTI for SOUL verification with Amali driving followed by LUCI ESM target for Barry

LBTI Verification overview:
  • (Y) Verifiy LBTI AO is delivering what is needed for Science run at end of May with M3 aligned on coldstop
  • Test updated 500 mode table (M3 aligned to WFS Moving in tip to -200, tilt 90)
  • (Y) Test 576 modes
LUCI ESM overview:
  • AO Ref star is R=11.2 (12.4 applying the Barry correction factor). This is Bin 1 ESM, however, it is ~2' from the dwarf galaxy.
  • Collecting "free" SL data with LUCI2
  • Master diagnostic patch has been activated
Doug powered up and checked out LUCI WFS in the afternoon and will log on for running the LUCI ESM show ~ 3:45UT

Overview:

2:14 Telescope opening up for business

2:15 Presetting to alpha Leo: V mag 1.40. This is for initial setup.

2:20 M3 LMIRcam position, global offset tip 70, tilt -30, absolute 590, 35.

2:23 Amali removed gross aberrations and is closing with 576 modes, 1700Hz.
Center pupils is hard coded to use modulation of 20 but you can't use modulation of 20 with a frequency of 1700Hz so center pupils fails

Updated ACE table with 200

EMg was 4 for mag 0.0-2.5 so updated that next since we were then getting OCAM illumination errors.

Wow, so no vignetting! What just happened. Warmer? Ellipse mirror position is good. So perhaps we want to move the WFS board to obtain more

Closed with 576 modes, freq 1700Hz, Estimated magnitude on the WFS 3.0 TN# 20210501_024429-024543 (pupil aligned on LMIRCAM at absolute tip 590, tilt 35)

Our unvignetted position in daytime the other day was an absolute position of: absolute tip 200 so moving to that to get closed loop data

The pupils do not look as flat as we would expect for 576 modes, some midrange being applied with optical gains. The winds are ~5m/s and we were seeing that amplitude ont eh adsec while pointing exactly out of the winds. Asked david to close the rear vent doors which reduced the amplitude of the winds on teh adsec to ~ 1m/s

3:12 DIMM 1.07"

Optimize gains is picking bad gains even though teh plots look lovely. The correction sems quite unstable. Could this be a 576 mode thing?
Taking LMIRCam and TN data for 576 mode with TN#030837 & and set of 6 TN#030927-031041. These are 576 mode 1700 freq, GOPT on, at old unvignetted tertiary position of absolute tip -200

Switching to 500 modes reconstructor., 1700Hz, Same tertiary pos, optimized. TN#031534-031702

Note about M3: Tertiary TIP movement moves down on CCD47 and left on lmircam

NOTE2: the 0.2 pupil offset threshold is in for LBTI too!

3:22 Moving to V=10.44 BD+ 16 2260

03:29 RIP when opening after center pupils. *

Amali recovered, flat set 638 modes.

Communication issues with adsec. The modal basis is not loading. Amali is resending.
typo in ace table, fixed and updated sim link. But we decided that we are going to pass on to LUCI now and will test the ACE table daytime.

3:54 Reauthorizong for LUCI

Main Outcome - We do not see the vignetting at any tertiary postion

Barry's ESM Program

04:00 David authorize as binoLUCI.

04:13 Pointed

04:27 TTM Preset failed because David had authorized at bentGregorianARGOS

04:33 TTM Preset failed. Out of focus on SX due to removal of ARGOS Dichroic

04:46 Finally pointed and collimated

04:46 TTM Preset
  • Close loop fail because camera lens did not converge. delta x 0.4

04:59 Closed loop
  • manually set gain to 3.5, 0.2
  • TN with some open loop 045917
  • TN 050023, 050456,051350, 051737
  • Barry says we're getting about a 33% improvement in FWHM, SX 0.4, DX 0.6

05;55 hand over to PEPSI

-- JenniferPower - 01 May 2021

-- DougMiller - 01 May 2021
Topic revision: r3 - 01 May 2021, DougMiller
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback